### v1

@@ 5 Conclusion
The work highlights the importance of balancing general and domain-specific abilities in LLMs and offers a promising approach to achieving this balance. Future research could explore broader applications of our block expansion method in other domains, for instance, it is an important task for multimodal large language models Ge et al. (2023); Bai et al. (2023) to preserve the original language ability.
Acknowledgements
We sincerely acknowledge Qingyue Zhao (Tsinghua University; ARC Lab, Tencent PCG) and Xiaohu Jiang (Tsinghua University; ARC Lab, Tencent PCG) for their engaging discussions.
@@ Appendix A Gradient Derivation
Appendix B MINT-Bench
@@
Table 9 displays the success rate for each model evaluated on various task type benchmarks, as well as the micro average when $k=5$. The LLaMA Pro - Instruct model demonstrates robust performance across all task types compared to other models of similar size. Figure  7 provides a case study to compare LLaMA Pro - Instruct and LLaMA2-7B-Chat where LLaMA Pro - Instruct successfully utilizes Python program to solve the given question in the multi-round interaction.
Appendix C MT-Bench
@@
MT-bench is a collection of demanding multi-turn open-ended questions designed for evaluating chat assistants. In order to automate the evaluation process, we employ powerful LLMs, such as GPT-4, to act as judges and assess the quality of the models’ responses. We present the detailed pairwise comparison in the Figure  8 and Figure  9. Figure  10 shows the case study of the comparison between LLaMA Pro - Instruct and LLaMA2-7B-Chat.
@@
Appendix D TRACE-Bench
Dataset Source Avg len Metric Language #data
Domain-specific
ScienceQA Science 210 Accuracy English 5,000
FOMC Finance 51 Accuracy English 5,000
MeetingBank Meeting 2853 ROUGE-L English 5,000
Multi-lingual
C-STANCE Social media 127 Accuracy Chinese 5,000
20Minuten News 382 SARI Germany 5,000
Code completion
Py150 Github 422 Edim similarity Python 5,000
Mathematical reasoning
NumGLUE-cm Math 32 Accuracy English 5,000
NumGLUE-ds Math 21 Accuracy English 5,000
Table 10: An overview of dataset statistics in TRACE. ’Source’ indicates the context’s origin. ’Avg len’ represents word count for English, German, and code datasets, and character count for Chinese. ’SARI’ is a score specific to simplification.
@@ Appendix D TRACE-Bench
C-STANCE 0.5 0.27 0.28 0.31 0.01 0.51 0.44 0.27
FOMC - 0.69 0.61 0.57 0.0 0.44 0.51 0.46
MeetingBank - - 0.472 0.287 0.159 0.111 0.14 0.139
Py150 - - - 0.61 0.294 0.551 0.547 0.364
ScienceQA - - - - 0.68 0.6 0.59 0.54
NumGLUE-cm - - - - - 0.395 0.309 0.272
NumGLUE-ds - - - - - - 0.59 0.51
20Minuten - - - - - - - 0.414
OP 0.371
BWT -0.173
Table 11: Detailed results of sequentially LoRA training of LLaMA2-7B.
[…]
C-STANCE 0.6 0.43 0.49 0.49 0.5 0.49 0.48 0.53
FOMC - 0.8 0.46 0.47 0.0 0.47 0.5 0.51
MeetingBank - - 0.523 0.338 0.344 0.392 0.396 0.371
Py150 - - - 0.594 0.265 0.561 0.566 0.575
ScienceQA - - - - 0.87 0.83 0.58 0.55
NumGLUE-cm - - - - - 0.383 0.198 0.210
NumGLUE-ds - - - - - - 0.64 0.49
20Minuten - - - - - - - 0.407
OP 0.455
BWT -0.147
Table 12: Detailed results of sequentially fine-tuning of LLaMA2-7B.
[…]
C-STANCE 0.65 0.52 0.52 0.52 0.28 0.52 0.5 0.57
FOMC - 0.8 0.49 0.45 0.0 0.44 0.46 0.45
MeetingBank - - 0.512 0.363 0.338 0.383 0.393 0.38
Py150 - - - 0.593 0.366 0.531 0.53 0.548
ScienceQA - - - - 0.89 0.8 0.6 0.63
NumGLUE-cm - - - - - 0.333 0.185 0.198
NumGLUE-ds - - - - - - 0.68 0.54
20Minuten - - - - - - - 0.408
OP 0.465
BWT -0.143
Table 13: Detailed results of sequentially block expansion training of LLaMA2-7B.
Table 10 shows the statistics of TRACE. We evaluate the performance of different training strategies after sequential tasks. Table 11-13 shows the detailed performance of different training strategies of each round during the continual learning with LLaMA2-7B.
Appendix E Domain of Law

### v2

@@ 6 Conclusion
7 Ethical Statement
LLaMA Pro and LLaMA Pro - Instruct are designed for a wide range of NLP tasks, with a focus on programming, mathematics, and general language tasks. It suits scenarios requiring integration of natural and programming languages. While LLaMA-Pro addresses some limitations of previous models in the series, it may still encounter challenges specific to highly specialized domains or tasks. Users should be aware of potential biases in the model and use it responsibly, considering its impact on various applications with the LLaMA-2 license.
@@ Appendix A Gradient Derivation
Appendix B Dataset Details
Data source Tokens Weight
Proof-Pile-2 55B 1.00
   AlgebraicStack 11B
   OpenWebMath 15B
   ArXiv 29B
The-Stack-Dedup
   Python 22B 1.50
ShareGPT User prompts GPT-3.5/GPT-4 63,817 2.9 293.2 1157.1
WizardLM_evol_instruct_V2 GPT-4 GPT-4 143,000 1.0 602.6 1704.9
SlimOrca Human-written GPT-4 517,982 1.0 574.3 599.3
MetaMath Human-written/GPT-4 GPT-4 395,000 1.0 209.4 498.2
Evol-CodeAlpaca GPT-4 GPT-4 111,272 1.0 652.5 1552.0
In this section, we provide detailed information about the dataset used for both pretraining and Supervised Fine-Tuning (SFT). Table 6 outlines the composition of our pretraining dataset, which comprises approximately 80 billion tokens from both math and code corpora. The specifics of the SFT data are delineated in Table 7.
For our proposed LLaMA Pro - Instruct, we employ a blend of multiple instruction datasets spanning general instruction, math, and code for the SFT process. These sources include ShareGPT [^3] https://huggingface.co/datasets/anon8231489123/ShareGPT_Vicuna_unfiltered, which contains real user and ChatGPT chat history records, and the WizardLM evolution instruction dataset Xu et al. (2023a), offering a wealth of instruction data with varying complexity levels. We also incorporate the evolution CodeAlpaca dataset Luo et al. (2023), which includes complex coding tasks generated by ChatGPT and their corresponding solutions. Additionally, we use MetaMath Yu et al. (2023), which reframes questions from multiple perspectives, and SlimOrca Lian et al. (2023), a curated subset of our OpenOrca data. SlimOrca provides an efficient route to achieve performance comparable to using larger data slices, while only incorporating approximately 500,000 GPT-4 completions.
Appendix C Mistal-Pro Details
Model GSM8k Pass@1 MATH Pass@1
MPT-7B 6.8 3.0
Falcon-7B 6.8 2.3
LLAMA-1-7B 11.0 2.9
LLAMA-2-7B 14.6 2.5
MPT-30B 15.2 3.1
LLAMA-1-13B 17.8 3.9
GPT-Neo-2.7B 19.5
Falcon-40B 19.6 2.5
Baichuan-chat-13B 23.9
Vicuna-v1.3-13B 27.6
LLAMA-2-13B 28.7 3.9
MetaMath-7B 66.5 19.8
MetaMath-13B 72.3 22.4
MetaMath-Mistral-7B 77.7 28.2
MetaMath-Llemma-7B 69.2 30.0
MetaMath-Mistral-Pro 78.4 30.3
Table 8: Performance of various models on GSM8k Pass@1 and MATH Pass@1
Mistral-Pro is an advanced version of the original Mistral model Jiang et al. (2023), enhanced through the addition of Transformer blocks. This version excels in combining general language understanding with domain-specific knowledge, particularly in programming and mathematics. It employs the same methodology for creating additional blocks as LLaMA-Pro but utilizes only $\frac{1}{10}$ of LLaMA Pro’s learning rate, as recommended by MetaMath-Mistral  [^4] https://huggingface.co/spaces/TencentARC/MetaMath-Mistral-Pro. We continued pretraining on code and math datasets, including the automath subset of Cosmopedia  [^5] https://huggingface.co/datasets/HuggingFaceTB/cosmopedia, proof-pile-2, and the Python subset of Stack. The supervised fine-tuning (SFT) approach remains consistent with MetaMath-Mistral, except that we switch the base model to our Mistral-Pro. The detailed results of GSM8k and MATH can be found in Table 8.
Appendix D Evaluation Benchmark
The benchmarks used for evaluation include:
Appendix E MINT-Bench
@@
Figure 9: The difference of three training strategies, fully fine-tuning, LoRA, and our proposed block expansion.
@@
Table 10 displays the success rate for each model evaluated on various task type benchmarks, as well as the micro average when $k=5$. The LLaMA Pro - Instruct model demonstrates robust performance across all task types compared to other models of similar size. Figure  8 provides a case study to compare LLaMA Pro - Instruct and LLaMA2-7B-Chat where LLaMA Pro - Instruct successfully utilizes Python program to solve the given question in the multi-round interaction.
Appendix F MT-Bench
@@
MT-bench is a collection of demanding multi-turn open-ended questions designed for evaluating chat assistants. In order to automate the evaluation process, we employ powerful LLMs, such as GPT-4, to act as judges and assess the quality of the models’ responses. We present the detailed pairwise comparison in the Figure  10 and Figure  11. Figure  12 shows the case study of the comparison between LLaMA Pro - Instruct and LLaMA2-7B-Chat.
@@
Appendix G Token Distribution
@@
lambada openai lambada standard stack
LLaMA-2-7B 3.39 4.13 9.46
LLaMA Pro 3.46 4.30 5.25
Table 11: The perplexity of LLaMA and LLaMA Pro evaluated across general domain and code domain.
We assess the token distribution between LLaMA-2-7B and LLaMA Pro, employing the methodology proposed by  Lin et al. (2023). Specifically, for a given user query $q=\{q_{1},q_{2},...\}$, we input it into LLaMA Pro to obtain its output $o=\{o_{1},o_{2},...\}$ using greedy decoding. For each position $t$, we define a context at this position as $x_{t}=q+\{o_{1},...,o_{t-1}\}$. We denote the aligned model’s probability distribution for predicting the next token at this position as $P_{\text{align}}$, where $o_{t}$ has the highest probability.
By passing the context $x_{t}$ into the base model LLaMA-2-7B, we generate another probability distribution, $P_{\text{base}}$, for sampling the next token at this position. First, the aligned model with greedy decoding is used to generate a full output $o$. For each position $t$, tokens are ranked according to their probability $P_{\text{base}}$ as predicted by the base model. The rank of $o_{t}$ in this sorted list is defined as the ’base rank’, denoted as $\eta$. This categorizes positions into three types: (1) unshifted positions ($\eta=1$): $o_{t}$ is the top-ranked token in both $P_{\text{base}}$ and $P_{\text{align}}$, having the highest probability; (2) marginal positions ($1<\eta\leq 3$): although $o_{t}$ is not the top-ranked token in $P_{\text{base}}$, it is still likely to be sampled for decoding, with the 2nd or 3rd highest probability; (3) shifted positions ($\eta>3$): in this case, $o_{t}$ is rather unlikely to be sampled by $P_{\text{base}}$, indicating a significant distribution shift from $P_{\text{base}}$ to $P_{\text{align}}$.
We conduct a perplexity evaluation of LLaMA-2-7B and LLaMA Pro across general and code corpora. For the general domain, we utilize two different versions of the LAMBADA dataset. For the code domain, we use the Python split of the bigcode/the-stack-smol-xs dataset [^6] https://huggingface.co/datasets/bigcode/the-stack-smol-xs. The results, presented in Table 11, indicate that LLaMA Pro effectively retains the language modeling ability for the general corpus while enhancing its proficiency in the code domain.
Appendix H Domain of Law
Figure 9 shows the difference between three training strategies that we use to conduct our ablation study. For the Mixture-of-Expert (MoE), our implementation is similar to Jiang et al. (2024). We use 2 experts and for each token, both experts will be activated. Specifically, We extend each FFN for all 32 layers, keep the original `$W_{3}`$ unchanged, learn an additional Linear layer with weights `$\hat{W_{3}}`$, and at the same time add two new learnable parameters `$\alpha_{1},\alpha_{2}`$ , when forward the output of Linear corresponding to $W_{3},\hat{W_{3}}$ will be weighted and summed with $\text{softmax}(\alpha_{1},\alpha_{2})$ and fed into the next block.